# Exercising the CCPA Opt-out Right on Android: Legally Mandated but Practically Challenging

### v5

@@ Abstract.
The business model of many mobile apps is based on generating revenue from sharing user data with ad networks and other companies to deliver personalized ads. The California Consumer Privacy Act (CCPA) gives California residents a right to opt out of the selling and sharing of their personal information. In two experiments we evaluate to which extent popular apps on the Android platform enable California residents to exercise their CCPA opt-out right. In our first experiment—manually exercising the opt-out right via app-level UIs for a set of 100 apps—we find that only 48 apps implement the legally mandated setting, which suggests that CCPA opt-out right non-compliance is a broader issue on the platform. In our second experiment—automatically exercising the opt-out right at the platform-level by sending Global Privacy Control (GPC) signals—we find for an app dataset of $1,811$ apps that GPC is largely ineffective. While we estimate with 95% confidence that 62%–81% of apps in our app dataset must respect the CCPA’s opt-out right, many apps do not do so. Disabling apps’ access to the AdID, which is not intended for exercising the CCPA opt-out right but could have practical effect in this regard, does not lead to a different result. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted in while only 26 had set it to opted out. Overall, our results suggest a compliance gap as California residents have no effective way of exercising their CCPA opt-out right on the Android platform; neither at the app- nor at the platform-level. We think that re-purposing the Android AdID setting as an opt-out right setting with legal meaning could resolve this compliance gap under the CCPA and other laws and improve users’ privacy on the platform overall.
global privacy control, gpc, opt out, app privacy, online privacy
@@ 1. Introduction
Most commonly mobile apps’ business model is based on advertising. For their revenue many apps rely on personalized advertising by sharing personal information with third-party ad networks, data brokers, and other companies in the online ad industry. However, such cross-context behavioral advertising is often not transparent and exposes users to privacy risks without control over the sharing of their data (Ullah et al., 2023). The California Consumer Privacy Act (CCPA) gives California residents a right to opt out of the selling and sharing of their personal information, including cross-context behavioral advertising. Mobile apps that are subject to the CCPA’s opt-out right are required to provide California residents functionality to exercise their right. Figure 1 shows an opt-out screen for one such app. In this study we evaluate to which extent popular apps on the Android platform enable California residents to exercise their CCPA opt-out right, how they can do so, and which effect, if any, exercising the right has. We focus on the Android platform as its apps have been shown to share personal information more widely than their iOS counterparts (Kollnig et al., 2022a).
@@
While app-level opt-out UIs, as shown in Figure 1 are one way to exercise the opt-out right, they are not the only one. The Office of the California Attorney General is also enforcing the opt-out right as expressed via Global Privacy Control (GPC) (State of California Department of Justice, 2022a, 2025), a technical specification for opting out (Zimmeck et al., 2025). Per the CCPA, a “business shall treat […] opt-out preference signal[s] as a valid request to opt-out of [the] sale/sharing […] for [a] browser or device. (California Privacy Protection Agency, 2024) [^1] CCPA Regulations, §7025(c)(1). Compliance with GPC is required in California since January 2021 (Becerra, 2021). However, achieving compliance is challenging for app operators since Android currently does not have a platform-level GPC setting. Thus, the effectiveness of GPC on Android is unclear and has not been studied in previous research. [^2] The same is true for iOS, other mobile platforms, IoT platforms, and, broadly, any other platform beyond the web. We see our study as a GPC readiness test of the Android platform: without platform-level support, to which extent do Android apps respect California residents’ opt out via GPC under the CCPA?
@@
Despite the lack of platform-level GPC support, there is already an opt-out setting on Android. While not intended for exercising legal opt-out rights, if a user has opted out using the AdID setting, Google prohibits app operators to use the AdID for creating user profiles for advertising purposes or for targeting users with personalized advertising (Google, 2024c). Google requires that all apps uploaded or published to Google Play must generally use the AdID in lieu of any other device identifier for any advertising purposes (Google, 2023). Thus, as part of our evaluation we also study the effect of disabling the AdID. We explore whether opting out using the AdID setting effectively results in exercising the CCPA opt-out right.
@@
    App-level UI Opt-out Evaluation: For a set of 100 popular Android apps, do those apps enable California residents to opt out per the CCPA, how can they do so, and which effect, if any, does exercising the opt-out right have? (Section 5)
@@
    Platform-level GPC Opt-out Evaluation: For a set of 1,811 popular apps, what effect does sending GPC signals have for opting out per the CCPA? What effect does disabling the AdID have? (Section 6)
By addressing these questions, we aim to advance legally mandated opt-out rights on mobile platforms. More broadly, we view our study as a contribution towards giving users more control over the sharing of their data on mobile platforms. The Android and iOS platforms are currently in a transitional phase. Mobile app operators are required to respect users’ GPC signals as valid means of exercising opt out rights in jurisdictions in which GPC is in effect. However, they lack platform-level support in form of native Android and iOS GPC settings and respective APIs. To that end, we proposes to re-purpose the current Android AdID setting as a platform-level opt-out setting with legal meaning (Section 7).
Addressing our research questions requires a combination of legal and technical analysis. In the following, we will discuss related work (Section 2), provide a brief legal evaluation of the CCPA opt-out right on mobile platforms (Section 3), and describe our methodology for detecting apps that are non-compliant with the CCPA’s opt out right (Section 4). [^3] All legal analyses in this study are performed by an author who is admitted to practice law in California and has multiple years of expertise working with regulatory agencies and companies on matters of California Privacy law. Based on this background we will then address our research questions and discuss the practical implications of our findings as we see them for the Android platform.
@@ 2.1. Dynamic Analysis of Mobile Apps’ Data Practices and Privacy Compliance
Our study makes use of dynamic analysis of mobile apps for purposes of privacy compliance. Early studies in this research area focused on modifying mobile operating systems to observe personal data traversing through the systems. Enck et al.’s TaintDroid (Enck et al., 2010) falls in this category for Android as does Agarwal’s and Hall’s ProtectMyPrivacy (Agarwal and Hall, 2013) for iOS. While this operating system-focused approach can give highly accurate insights into apps’ data flows, the growing complexity of mobile operating systems makes it increasingly challenging. Thus, more recent studies tend to focus on the capture and analysis of apps’ network traffic (Ren et al., 2016; Song and Hengartner, 2015; Le et al., 2015; Reardon et al., 2019), which is also our approach. In the closest work to ours, Kollnig et al. performed a dynamic analysis of the network traffic of 1,297 Android apps and found that while most engage in third-party tracking, only few obtained user consent before doing so (Kollnig et al., 2021). While they selected their apps from the Google Play Store randomly, we selected the apps for our study based on Google’s lists of popular apps as those are most likely to rely on an ad-based business model that implies selling personal information or sharing such with third parties (Section 3.1). Nguyen et al. (Nguyen et al., 2021, 2022) extended the work of Kollnig et al. and found that 24,838/86,163 (29%) apps sent personal data to data controllers without the user’s prior explicit consent. While previous studies focused on EU law, we focus on the CCPA.
In many cases, the lack of consent is the result of a coordination failure. Specifically, most third parties rely on app operators to obtain consent on their behalf but in many cases fail to put in place robust measures to ensure operators can actually do so (Kollnig et al., 2021). Consent Management Platforms (CMPs), such as OneTrust (OneTrust, 2024), provide software integrations to assist operators in obtaining consent. For Apple devices, the platform itself includes a consent setting. With iOS 14.5 Apple introduced mandatory choice prompts. In doing so, Apple deviated from its earlier implementation of the ad tracking setting as an opt-out that used to be similar to that implemented on Android since 2013. However, Apple is facing challenges to enforce its App Tracking Transparency framework (Kollnig et al., 2022b). Independently from the platform, there are also usability challenges. Examining 100,000 Android apps, Chen et al. found that 36.3% of privacy settings identified in these apps are hard to find due to unclear user interface designs, which impact users’ privacy because 82.2% of the observed settings are set to disclose data by default (Chen et al., 2019).
@@ 2.2. Opting Out of Ad Tracking in Mobile Apps
Users often do not know which data is collected from them and by whom (Zimmeck et al., 2024a). Industry concentration has significantly increased the breadth and depth of data accessible to individual companies engaging in data collection (Binns et al., 2018). For a long time mobile platforms have had dedicated identifiers to enable third-party tracking. On iOS, Apple rolled out the Identifier for Advertisers, alongside an opt-out setting from ad tracking, in 2012 (Testwuide, 2022). Google quickly followed suit for Android by rolling out similar features in 2013 (Testwuide, 2022). These initiatives had the goal of moving away from using permanent device identifiers, such as the Unique Device Identifier on iOS or the Android_ID on Android, for advertising purposes. Permanent identifiers cannot be reset by users and, thus, present an obstacle for opting out. In the wake of these developments, various self-help tools for blocking ad tracking have emerged (Le et al., 2015; Greene and Shilton, 2018; Kollnig and Shadbolt, 2022).
To exercise their legal opt-out rights for various US jurisdictions, users can use Consumer Reports’ Permission Slip app, which has a human in the loop to perform the actual opt-out (Consumer Reports, 2024). Further, OptOutCode is a platform-agnostic proposal to automate opting out based on users prefixing their devices’ names with “0\$S” for “do not (0) sell (\$) or share (S)” personal information (Privacy4Cars, 2024). [^4] This idea is similar to Google’s opt-out solution for Google Location services access points (Google Maps Help, 2024): “To opt out, change the SSID (name) of your Wi-Fi access point (your wireless network name) so that it ends with ‘_nomap.’ For example, if your SSID is ‘12345,’ you would change it to ‘12345_nomap.”’ The Digital Advertising Alliance, an ad industry association, provides the AppChoices app (Digital Advertising Alliance, 2024). However, despite these tools, compared to websites, the choices for users to exercise their opt-out rights have traditionally been more limited on mobile apps. This development is the result of the lag of privacy feature implementations on mobile platforms.
@@ 2.3. Privacy Preference Signals
Recently, privacy preference signals have re-emerged as a viable opt-out method (Hils et al., 2021). However, various legal, technical, usability, and accountability challenges persist (Human et al., 2022). The Platform for Privacy Preferences Project (P3P) was an early signal to help people understand data practices and automate opt-outs and other privacy choices based on machine-readable privacy policies (Cranor et al., 2002, 2006). Advanced Data Protection Control (ADPC) follows in P3P’s footsteps and is developed for EU law (Advanced Data Protection Control (2022), ADPC). Perhaps, the most well-known signal, Do Not Track (DNT) (Fielding and Singer, 2019), was developed from 2009 as a binary signal for users to express their opt-out of tracking towards websites and other online services (California Legislative Information, 2003). However, DNT was never broadly adopted as there is no law mandating compliance. Recently, a decision of the Landgericht Berlin (Landgericht Berlin, 2023)—requiring LinkedIn to respect DNT signals—revitalized DNT. It remains to be seen whether this case remains an outlier. As a successor to DNT, GPC (Zimmeck et al., 2025) is currently emerging as a standard for exercising opt-out rights.
In this study, we are interested in GPC (Zimmeck et al., 2025). GPC is a binary signal with which users can express their Do Not Sell Or Share preference (Zimmeck et al., 2025). Most importantly, conforming to this preference is required under the CCPA and has been enforced by the Office of the California Attorney General (State of California Department of Justice, 2022a, 2025). Our GPC analysis is based on the methodology developed by Zimmeck et al. for the web (Zimmeck et al., 2023), which we adapt for use on the Android platform. The basic idea is to inject GPC headers into all apps’ outgoing HTTP requests to detect changes in app behavior that could indicate whether the opt-out via GPC is respected or not. For example, we evaluate whether GPC signals are propagated to third parties by measuring respective changes in their US Privacy Strings (IAB, 2024). If apps do not notify ad networks and other third parties of California residents’ opt-out preferences via privacy flags, such omission indicates a CCPA violation (Zimmeck et al., 2023). We also measure the effect of disabling the AdID.
@@ 3. Legal Background
3.1. The CCPA’s Opt-out Right
The CCPA gives consumers, i.e., California residents, [^5] CCPA, §1798.140(i). a right to opt out. [^6] CCPA, §1798.120. The right does not apply to residents of other states or countries. Specifically, a consumer has the right “at any time, to direct a business that sells or shares personal information about the consumer to third parties not to sell or share the consumer’s personal information.” [^7] CCPA, §1798.120(a) (emphasis added). Whether an app operator is running a “business” depends on annually buying, selling, or sharing the personal information of 100,000 or more consumers or households. [^8] Alternatively, thresholds of annual gross revenues in excess of 25 million dollars or deriving 50% or more of annual revenues from selling or sharing consumers’ personal information can be used to determine whether an entity qualifies as a “business” per CCPA, §1798.140(d)(1). For this determination Samarin et al. used the number of app installs (Samarin et al., 2023). We rely on app operators’ self-declaration of CCPA applicability in their privacy policies (Section 4.4), which represents a lower bound since there may be additional apps that are subject to CCPA but do not declare that to be the case. This approach also has the advantage of identifying a business across apps. For example, an operator may have multiple apps, each of which does not reach the threshold for CCPA applicability on its own.
For the opt-out right to be applicable, the app operator must also “sell” or “share” personal information. “Selling” means to communicate a consumer’s personal information to a third party for monetary or other valuable consideration. [^9] CCPA, §1798.140(ad)(1). “Sharing” means to communicate a consumer’s personal information to a third party for cross-context behavioral advertising. [^10] CCPA, §1798.140(ah)(1). The concept of “sharing” was introduced to the CCPA by the California Privacy Rights Act, an amendment to the CCPA, to remove any doubt that the CCPA applies to cross-context behavioral advertising (TrueVault, 2022). Evidence for selling or sharing can consist of dedicated CCPA opt-out of sale or share links, respective opt-out implementations, or statements of applicability of the CCPA opt-out right in privacy policies (O’Connor et al., 2020). As the enforcement actions of the Office of the California Attorney General show (State of California Department of Justice, 2025), the integration of third-party libraries for ads personalization falls squarely under the definitions of selling and sharing. Thus, for our GPC opt-out evaluation we rely on either the app operator disclosing in its privacy policy that it is selling or sharing personal information or the integration of an ad network that discloses in its privacy policy that it is “buying” or “collecting” personal information or “selling” or “sharing” such downstream to other companies.

### v6

@@ Abstract.
Many mobile apps’ business model is based on sharing user data with ad networks to deliver personalized ads. The California Consumer Privacy Act (CCPA) gives California residents a right to opt out. In two experiments we evaluate to which extent popular Android apps enable California residents to exercise their right. In our first experiment—manually exercising the right via app-level UIs—we find that only 48 out of 100 apps implement a respective setting, which suggests that CCPA opt-out right compliance on the Android platform is generally low. In our second experiment—automatically exercising the opt-out right by sending Global Privacy Control (GPC) signals—we find for an app dataset of $1,811$ apps that GPC is largely ineffective. While we estimate with 95% confidence that 62%–81% of apps in our app dataset must respect the CCPA opt-out right, many apps do not do so. Our evaluation of disabling apps’ access to the AdID—which is technically not intended for exercising the CCPA opt-out right but could be practically effective—does not change our conclusion. For example, when sending GPC signals and disabling apps’ access to the AdID, 338 apps still had the ccpa status of the ad network Vungle set to opted_in while only 26 had set it to opted_out. Overall, our results suggest a compliance gap as California residents have no effective way of exercising their CCPA opt-out right on the Android platform; neither at the app- nor at the platform-level. We think that re-purposing the Android AdID setting as an opt-out right setting with legal meaning under the CCPA and other laws could close this gap and improve users’ privacy on the platform significantly.
global privacy control, gpc, opt-out, app privacy, online privacy
@@ 1. Introduction
The business model of many mobile apps is based on advertising. Many apps rely on personalized advertising and share personal information with third-party ad networks, data brokers, and other companies in the online ad industry. However, such cross-context behavioral advertising is often not transparent and exposes users to privacy risks without control over the sharing of their data (Ullah et al., 2023). The California Consumer Privacy Act (CCPA) gives California residents a right to opt out of the selling and sharing of their personal information, including cross-context behavioral advertising. Mobile apps that are subject to the CCPA opt-out right are required to provide California residents functionality to exercise their right. Figure 1 shows an opt-out screen for one such app. In this study we evaluate to which extent popular apps on the Android platform enable California residents to exercise their CCPA opt-out right, how they can do so, and which effect, if any, exercising the right has. We focus on the Android platform as its apps have been shown to share personal information more broadly than their counterparts on the iOS platform (Kollnig et al., 2022a).
@@
While app-level opt-out UIs, as shown in Figure 1, are one way to exercise the opt-out right, they are not the only way. The Office of the California Attorney General is also enforcing the opt-out right as expressed via Global Privacy Control (GPC) (State of California Department of Justice, 2022a, 2025), a technical specification for opting out (Zimmeck et al., 2025). Per the CCPA, a “business shall treat […] opt-out preference signal[s] as a valid request to opt-out of [the] sale/sharing […] for [a] browser or device. (California Privacy Protection Agency, 2024) [^1] CCPA Regulations §7025(c)(1). Compliance with GPC is required in California since January 2021 (Becerra, 2021). However, achieving compliance is challenging for app operators since Android currently does not have a platform-level GPC setting. [^2] The same is true for iOS, other mobile platforms, IoT platforms, and, broadly, any other platform beyond the web. Thus, the effectiveness of GPC on Android is unclear and has not been studied in previous research. We see our study as a GPC requirement evaluation of the Android platform: without platform-level GPC support, to which extent do Android apps respect California residents’ opt-out rights under the CCPA?
@@
Despite the lack of platform-level GPC support, there is already an opt-out setting on Android. While not intended for exercising legal opt-out rights, if a user has opted out using the AdID setting, Google prohibits app operators to use the AdID for creating user profiles for advertising purposes or for targeting users with personalized advertising (Google, 2025e). Google requires that all apps uploaded or published to the Google Play store must generally use the AdID in lieu of any other device identifier for any advertising purposes (Google, 2025a). Thus, as part of our evaluation we also study the effect of disabling the AdID. We explore whether opting out using the AdID setting effectively results in exercising the CCPA opt-out right.
@@
    UI Opt-out Evaluation: For a set of 100 popular Android apps, do those apps enable California residents to opt out per the CCPA, how can they do so, and which effect, if any, does exercising the opt-out right have? (Section 5)
@@
    GPC Opt-out Evaluation: For a set of 1,811 popular Android apps, what effect does sending GPC signals have for opting out per the CCPA? What effect does disabling the AdID have? (Section 6)
By addressing these questions, we aim to advance legally mandated opt-out rights on mobile platforms. More broadly, we view our study as a contribution towards giving users more control over the sharing of their data on mobile platforms. The Android and iOS platforms are currently in a transitional phase. Mobile app operators are required to respect users’ GPC signals as valid means of exercising opt-out rights in jurisdictions in which GPC is in effect. However, they lack platform-level support in form of native Android and iOS GPC settings and respective APIs. To that end, we proposes to re-purpose the current Android AdID setting as a platform-level opt-out setting with legal meaning (Section 7).
Addressing our research questions requires a combination of legal and technical analysis. In the following, we will discuss related work (Section 2), provide a brief background discussion on the application of the CCPA opt-out right on mobile platforms (Section 3), and describe our methodology for detecting apps that are non-compliant with the right (Section 4). Based on this groundwork we will then address our research questions and discuss the practical implications of our findings for the Android platform.
@@ 2.1. Dynamic Analysis of Mobile Apps’ Data Practices and Privacy Compliance
Our study makes use of dynamic analysis of mobile apps for purposes of privacy compliance analysis. Early studies in this research area focused on modifying mobile operating systems to observe personal data traversing through the systems. Enck et al.’s TaintDroid (Enck et al., 2010) falls in this category for Android as does Agarwal’s and Hall’s ProtectMyPrivacy (Agarwal and Hall, 2013) for iOS. While this operating system-focused approach can give highly accurate insights into apps’ data flows, the growing complexity of mobile operating systems makes it increasingly challenging. Thus, more recent studies tend to focus on the capture and analysis of apps’ network traffic (Ren et al., 2016; Song and Hengartner, 2015; Le et al., 2015; Reardon et al., 2019), which is also our approach. In the closest work to ours, Kollnig et al. performed a dynamic analysis of the network traffic of 1,297 Android apps and found that while most engaged in third-party tracking, only few obtained user consent before doing so (Kollnig et al., 2021). While they selected their apps from the Google Play store randomly, we selected apps based on Google’s lists of popular apps as those are most likely to rely on an ad-based business model that implies selling of personal information or sharing such with third parties (Section 3.1). Nguyen et al. (Nguyen et al., 2021, 2022) extended the work of Kollnig et al. and found that 24,838/86,163 (29%) apps sent personal data to data controllers without the user’s prior explicit consent. While previous studies focused on EU law, we focus on the CCPA.
@@ 2.2. Opting Out of Ad Tracking in Mobile Apps
In many cases, the lack of consent is the result of a coordination failure. Specifically, most third parties rely on app operators to obtain consent on their behalf but in many cases fail to put in place robust measures to ensure operators can actually do so (Kollnig et al., 2021). Consent Management Platforms (CMPs), such as OneTrust (OneTrust, 2024), provide software integrations to assist operators in obtaining consent. For Apple devices, the platform itself includes a consent setting. With iOS 14.5 Apple introduced mandatory choice prompts. In doing so, Apple deviated from its earlier implementation of the ad tracking setting as an opt-out mechanism similar to that implemented on Android since 2013. However, Apple is facing challenges to enforce its App Tracking Transparency framework (Kollnig et al., 2022b). Across platforms, there are also usability challenges. Examining 100,000 Android apps, Chen et al. found that 36.3% of privacy settings identified in these apps are hard to find due to unclear user interface designs, which impact users’ privacy because 82.2% of the observed settings are set to disclose data by default (Chen et al., 2019).
Users often do not know which data is collected from them and by whom (Zimmeck et al., 2024a). Industry concentration has significantly increased the breadth and depth of data accessible to individual companies engaging in data collection (Binns et al., 2018). For a long time mobile platforms have had dedicated identifiers to enable third-party tracking. On iOS, Apple rolled out the Identifier for Advertisers, alongside an opt-out setting from ad tracking, in 2012 (Testwuide, 2022). Google quickly followed suit for Android by rolling out a similar mechanism in 2013 (Testwuide, 2022). These initiatives had the goal of moving away from using permanent device identifiers, such as the Unique Device Identifier on iOS or the Android_ID on Android, for advertising purposes. Permanent identifiers cannot be reset by users and, thus, entail identification risk. In the wake of these developments, various self-help tools for blocking ad tracking have emerged (Le et al., 2015; Greene and Shilton, 2018; Kollnig and Shadbolt, 2022).
To exercise their legal opt-out rights for various US jurisdictions, users can use Consumer Reports’ Permission Slip app, which has a human in the loop for performing some opt-out tasks (Consumer Reports, 2025). Further, OptOutCode is a platform-agnostic proposal to automate opting out based on users’ prefixing the names of their devices with “0\$S” for “do not (0) sell (\$) or share (S)” personal information (Privacy4Cars, 2025). [^3] This idea is similar to Google’s opt-out solution for Google Location services access points (Google Maps Help, 2025): “To opt out, change the SSID (name) of your Wi-Fi access point (your wireless network name) so that it ends with ‘_nomap.’ For example, if your SSID is ‘12345,’ you would change it to ‘12345_nomap.”’ The Digital Advertising Alliance, an ad industry association, provides the AppChoices app (Digital Advertising Alliance, 2025). However, despite these tools, privacy choices have historically been more limited on mobile platforms compared to the web. This development is the result of the lag of privacy mechanism implementations on mobile platforms.
@@ 2.3. Privacy Preference Signals
Recently, privacy preference signals have re-emerged and present an increasingly viable opt-out method (Hils et al., 2021). However, various legal, technical, usability, and accountability challenges persist (Human et al., 2022). Automating privacy choices via privacy preference signals has been tried before. The Platform for Privacy Preferences Project (P3P) was an early signal to help people understand data practices and automate opt-outs and other privacy choices based on machine-readable privacy policies (Cranor et al., 2002, 2006). Advanced Data Protection Control (ADPC) follows in P3P’s footsteps and is developed for EU law (Human et al., 2021). Perhaps, the most well-known signal, Do Not Track (DNT) (Fielding and Singer, 2019), was developed from 2009 as a binary signal for users to express their opt-out of tracking towards websites and other online services (California Legislative Information, 2003). However, DNT, just as the other signals, was never broadly adopted because no law mandates compliance. Recently, a decision of the Landgericht Berlin (Landgericht Berlin, 2023)—requiring LinkedIn to respect DNT signals—revitalized DNT. However, so far the decision remains an outlier. As a successor to DNT, GPC (Zimmeck et al., 2025) is currently emerging as the standard for exercising opt-out rights. Thus, in our study we focus on GPC.
GPC is a binary signal with which users can express their Do Not Sell Or Share preference (Zimmeck et al., 2025). Most importantly, compliance with this preference is required under the CCPA and has been enforced by the Office of the California Attorney General (State of California Department of Justice, 2022a, 2025). Our GPC analysis is based on the methodology developed by Zimmeck et al. for the web (Zimmeck et al., 2023), which we adapt for use on the Android platform. The basic idea is to inject GPC headers into all apps’ outgoing HTTP requests to detect changes in app behavior that could indicate whether the opt-out via GPC is respected or not. For example, we evaluate whether GPC signals are propagated to third parties by measuring respective changes in their US Privacy Strings (IAB Tech Lab, 2025). If apps do not notify ad networks and other third parties of California residents’ opt-out preferences by setting privacy flags, such omission indicates a CCPA violation (Zimmeck et al., 2023). In addition to measuring the effect of sending GPC signals, we also measure the effect of disabling the AdID.
@@ 3. Legal Background
3.1. The CCPA Opt-out Right
The CCPA gives consumers, i.e., California residents, [^4] CCPA §1798.140(i). a right to opt out. [^5] CCPA §1798.120. The right does not apply to residents of other states or countries. Specifically, a consumer has the right “at any time, to direct a business that sells or shares personal information about the consumer to third parties not to sell or share the consumer’s personal information.” [^6] CCPA §1798.120(a) (emphasis added). Whether an app operator is running a “business” depends on annually buying, selling, or sharing the personal information of 100,000 or more consumers or households. [^7] Alternatively, thresholds of annual gross revenues in excess of 25 million dollars or deriving 50% or more of annual revenues from selling or sharing consumers’ personal information can be used to determine whether an entity qualifies as a “business” per CCPA §1798.140(d)(1). For this determination Samarin et al. used the number of app installs (Samarin et al., 2023). We rely on app operators’ self-declaration of CCPA applicability in their privacy policies (Sections 4.3 and 4.4), which represents a lower bound since there may be additional apps that are subject to the CCPA but whose operators do not declare so. This approach also has the advantage of identifying a business across apps. For example, an operator may have multiple apps, each of which does not reach the threshold for CCPA applicability on its own.
For the opt-out right to be applicable, the app operator must also “sell” or “share” personal information. “Selling” means to communicate a consumer’s personal information to a third party for monetary or other valuable consideration. [^8] CCPA §1798.140(ad)(1). “Sharing” means to communicate a consumer’s personal information to a third party for cross-context behavioral advertising. [^9] CCPA §1798.140(ah)(1). The concept of “sharing” was introduced to the CCPA by the California Privacy Rights Act, an amendment to the CCPA, to remove any doubt that the CCPA applies to cross-context behavioral advertising (TrueVault, 2025). Evidence for selling or sharing can consist of dedicated CCPA opt-out of sale or share links, respective opt-out implementations, or statements of applicability of the CCPA opt-out right in privacy policies (O’Connor et al., 2021). As the enforcement actions of the Office of the California Attorney General show (State of California Department of Justice, 2025), the integration of third-party libraries for ads personalization falls squarely under the definitions of selling and sharing. Thus, for our GPC opt-out evaluation we rely on an app’s integration of an ad network that discloses in its privacy policy that it is “buying” or “collecting” personal information or “selling” or “sharing” such (Section 4.4).